# On the Mean Value of a Weighted Composite Arithmetic Function

### v2

@@ 1 Introduction
$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right),$
@@
$\mathcal{C}(r,k)=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
@@
$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
@@ 2 Preliminary Lemmas
$\sum_{n\leq X}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{X^{s}}{s}ds+O\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{X}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(X/n)|}\right)\right).$
@@ 3 Proof of the Main Result
This explicit form allows us to isolate the dominant poles. Through some manipulation and comparison with Euler’s products of the Riemann zeta function, we can factor the general series as follows
@@
where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$. Specifically, $H(s)$ contains the remaining convergent factors
This factorization confirms that $f(s)$ possesses a pole of order 2 at $s=1$, arising from the $\zeta(s)^{2}$ term.
For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula allows us to express the summatory function as
@@
Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $O(x^{\epsilon})$.
@@
$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+O(h^{2})\right),$
@@
$\frac{x^{s}}{s}=x\left(1+h\ln x+O(h^{2})\right)\left(1-h+O(h^{2})\right)\\ =x\left(1+h(\ln x-1)+O(h^{2})\right).$
@@
$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+O\left((s-1)^{2}\right).$
@@
$f(s)=\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1),\quad\text{as }s\to 1.$
where $A$ and $B$ are constants determined by the Euler product of $f(s)$.
@@
$f(s)\frac{x^{s}}{s}=\left(\frac{A}{(s-1)^{2}}+\frac{B}{s-1}+O(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+O\left((s-1)^{2}\right)\right)$
$\frac{Ax}{(s-1)^{2}}+\frac{Ax(\ln x-1)}{s-1}+\frac{Bx}{s-1}+O(1).$
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax(\ln x-1)+Bx=Ax\ln x+(B-A)x.$
@@
$H(s)=\frac{\zeta(ks)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right).$
The constant $A$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
$A=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$
@@
$A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right).$
This constant $A$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
The constant $B$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
$\zeta(s)=\frac{1}{s-1}+\gamma+O(s-1),$
@@
$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+O(1).$
Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+O((s-1)^{2})$ and comparing coefficients yields
$B=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma A.$
@@
$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(k)}{\zeta(k)}\cdot k-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{ks}+p^{(k-1)s})}\right)\right]_{s=1}.$
Each term in this expression is explicitly computable: $\zeta(k)$ and $\zeta^{\prime}(k)$ are known values for integer $k\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=Ax\ln x+(B-A)x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
@@
$\mathcal{C}(r,k)=A=\frac{6\zeta(k)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{k}+p^{k-1})}\right),$
@@
$\mathcal{K}=B-A=H^{\prime}(1)+(2\gamma-1)A.$
It remains to bound the integrals along the new segments of the contour to establish the error term $O(x^{\frac{1}{2}+\epsilon})$.
On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
With $T=x$, this contribution is typically bounded by $O(x^{\frac{1}{2}+\epsilon})$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
@@
The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll x^{\frac{1}{2}+\epsilon}.$
@@
$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+O\left(x^{\frac{1}{2}+\epsilon}\right).$
@@
$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+O\left(x^{\frac{1}{2}+\epsilon}\right).$

### v3

@@ 1 Introduction
$\sum_{n\leq x}D_{k}(\mathcal{M}_{r}(n))=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+\mathcal{O}\left(x^{\frac{1}{2}+\epsilon}\right),$
@@
$\mathcal{C}(r,k)=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right),$
@@ 2 Preliminary Lemmas
$\sum_{n\leq x}a_{n}=\frac{1}{2\pi i}\int_{b-iT}^{b+iT}F(s)\frac{x^{s}}{s}ds+\mathcal{O}\left(\sum_{n=1}^{\infty}|a_{n}|\left(\frac{x}{n}\right)^{b}\min\left(1,\frac{1}{T|\ln(x/n)|}\right)\right).$
@@ 3 Proof of the Main Result
To facilitate the factorization, we combine the terms over a common denominator
$\text{L.F}=\frac{k(1-p^{-s})(1-p^{-rs})+p^{-s}(2-p^{-rs})}{k(1-p^{-s})(1-p^{-rs})}\\$
$\frac{k(1-p^{-s}-p^{-rs}+p^{-(r+1)s})+2p^{-s}-p^{-(r+1)s}}{k(1-p^{-s})(1-p^{-rs})}\\$
$\frac{k+(2-k)p^{-s}-kp^{-rs}+(k-1)p^{-(r+1)s}}{k(1-p^{-s})(1-p^{-rs})}.$
This explicit rational form allows us to isolate the dominant poles by comparing with the Euler products of the Riemann zeta-function.
Recall that the Riemann zeta-function has the Euler product $\zeta(s)=\prod_{p}(1-p^{-s})^{-1}$. To obtain a pole of order 2 at $s=1$ (consistent with the $x\ln x$ term in the asymptotic formula), we extract the factor $\zeta(s)^{2}=\prod_{p}(1-p^{-s})^{-2}$.
We rewrite the local factor as
$\frac{1}{(1-p^{-s})^{2}}\cdot\underbrace{\frac{(1-p^{-s})\left[k+(2-k)p^{-s}-kp^{-rs}+(k-1)p^{-(r+1)s}\right]}{k(1-p^{-rs})}}_{H_{p}(s)}.$
Taking the product over all primes, we obtain the global factorization
@@
where $H(s)=\prod_{p}H_{p}(s)$. To express $H(s)$ in terms of standard zeta-functions, we note that the denominator contains $(1-p^{-rs})$, which corresponds to $\zeta(rs)=\prod_{p}(1-p^{-rs})^{-1}$. After careful manipulation and normalization (ensuring $H(1)$ converges to a non-zero constant), we obtain
$H(s)=\frac{\zeta(rs)}{\zeta(2s)}\prod_{p}\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right),$
where $H(s)$ is analytic and non-zero in the region $\text{Re}(s)\geq\frac{1}{2}+\epsilon$.
For $x\geq 1$, $b>1$ (the abscissa of absolute convergence), and $T>0$, Perron’s formula (see Lemma 4) allows us to express the summatory function as
@@
Under the assumption of polynomial growth $|F_{r,k}(n)|\ll n^{\epsilon}$, and by choosing $b=1+\epsilon$ and $T=x$, the error term becomes negligible compared to the main term, typically of the order $\mathcal{O}(x^{\epsilon})$.
@@
$x^{s}=x^{1+h}=x\cdot e^{h\ln x}=x\left(1+h\ln x+\mathcal{O}(h^{2})\right),$
@@
$\frac{x^{s}}{s}=x\left(1+h\ln x+\mathcal{O}(h^{2})\right)\left(1-h+\mathcal{O}(h^{2})\right)\\ =x\left(1+h(\ln x-1)+\mathcal{O}(h^{2})\right).$
@@
$\frac{x^{s}}{s}=x\left(1+(s-1)(\ln x-1)\right)+\mathcal{O}\left((s-1)^{2}\right).$
@@
$f(s)=\frac{\mathcal{A}}{(s-1)^{2}}+\frac{\mathcal{B}}{s-1}+\mathcal{O}(1),\quad\text{as }s\to 1.$
where $\mathcal{A}$ and $\mathcal{B}$ are constants determined by the Euler product of $f(s)$.
@@
$f(s)\frac{x^{s}}{s}=\left(\frac{\mathcal{A}}{(s-1)^{2}}+\frac{\mathcal{B}}{s-1}+\mathcal{O}(1)\right)\cdot\left(x+x(\ln x-1)(s-1)+\mathcal{O}\left((s-1)^{2}\right)\right)$
$\frac{\mathcal{A}x}{(s-1)^{2}}+\frac{\mathcal{A}x(\ln x-1)}{s-1}+\frac{\mathcal{B}x}{s-1}+\mathcal{O}(1).$
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=\mathcal{A}x(\ln x-1)+\mathcal{B}x=\mathcal{A}x\ln x+(\mathcal{B}-\mathcal{A})x.$
@@
The constant $\mathcal{A}$ corresponds to the residue of the double pole and is determined by evaluating $H(s)$ at $s=1$,
$\mathcal{A}=\lim_{s\to 1}(s-1)^{2}f(s)=H(1).$
@@
$\mathcal{A}=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right).$
This constant $\mathcal{A}$ corresponds precisely to the leading coefficient $\mathcal{C}(r,k)$ in our main theorem.
The constant $\mathcal{B}$ arises from the simple pole term and requires the derivative $H^{\prime}(1)$. Using the Laurent expansion of $\zeta(s)$ near $s=1$,
$\zeta(s)=\frac{1}{s-1}+\gamma+\mathcal{O}(s-1),$
@@
$\zeta(s)^{2}=\frac{1}{(s-1)^{2}}+\frac{2\gamma}{s-1}+\mathcal{O}(1).$
Multiplying by $H(s)=H(1)+H^{\prime}(1)(s-1)+\mathcal{O}((s-1)^{2})$ and comparing coefficients yields
$\mathcal{B}=H^{\prime}(1)+2\gamma H(1)=H^{\prime}(1)+2\gamma\mathcal{A}.$
@@
$\frac{H^{\prime}(1)}{H(1)}=\frac{\zeta^{\prime}(r)}{\zeta(r)}\cdot r-2\frac{\zeta^{\prime}(2)}{\zeta(2)}+\sum_{p}\frac{d}{ds}\left[\ln\left(1-\frac{1}{k(p^{rs}+p^{(r-1)s})}\right)\right]_{s=1}.$
Each term in this expression is explicitly computable: $\zeta(r)$ and $\zeta^{\prime}(r)$ are known values for integer $r\geq 2$, $\zeta(2)=\frac{\pi^{2}}{6}$, $\zeta^{\prime}(2)=\frac{\pi^{2}}{6}(\gamma+\ln(2\pi)-12\ln\lambda)$ where $\lambda\approx 1.2824$ is the Glaisher-Kinkelin constant, and the sum over primes converges rapidly.
@@
$\text{Res}_{s=1}\left(f(s)\frac{x^{s}}{s}\right)=\mathcal{A}x\ln x+(\mathcal{B}-\mathcal{A})x=\mathcal{C}(r,k)x\ln x+\mathcal{K}x,$
@@
$\mathcal{C}(r,k)=\mathcal{A}=\frac{6\zeta(r)}{\pi^{2}}\prod_{p}\left(1-\frac{1}{k(p^{r}+p^{r-1})}\right),$
@@
$\mathcal{K}=\mathcal{B}-\mathcal{A}=H^{\prime}(1)+(2\gamma-1)\mathcal{A}.$
It remains to bound the integrals along the new segments of the contour to establish the error term $\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right)$.
On the segments from $b\pm iT$ to $\sigma_{0}\pm iT$, we use convexity bounds for the zeta-function provided by Lemma 5 (e.g., $|\zeta(\sigma+iT)|\ll T^{\mu(\sigma)}$)
$\left|\int_{\sigma_{0}}^{b}f(\sigma\pm iT)\frac{x^{\sigma\pm iT}}{\sigma\pm iT}\,d\sigma\right|\ll_{r,k}\frac{x^{b}}{T}\int_{\sigma_{0}}^{b}|f(\sigma\pm iT)|\,d\sigma.$
With $T=x$, this contribution is typically bounded by $\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right)$. On the line $\text{Re}(s)=\sigma_{0}=\frac{1}{2}+\epsilon$, the integral is
@@
The magnitude is dominated by $x^{\sigma_{0}}=x^{\frac{1}{2}+\epsilon}$. Using Lemma 5 and the mean value theorem for the zeta-function (integral of $|\zeta(\frac{1}{2}+it)|^{2}$), we obtain
$\int_{-T}^{T}\left|f\left(\frac{1}{2}+\epsilon+it\right)\right|\frac{x^{\frac{1}{2}+\epsilon}}{|t|}\,dt\ll_{r,k}x^{\frac{1}{2}+\epsilon}\cdot T^{\epsilon}\ll_{r,k}x^{\frac{1}{2}+\epsilon}.$
@@
$\frac{1}{2\pi i}\int_{b-iT}^{b+iT}f(s)\frac{x^{s}}{s}\,ds=\text{Res}_{s=1}+\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right).$
@@
$\sum_{n\leq x}F_{r,k}(n)=\mathcal{C}(r,k)x\ln x+\mathcal{K}x+\mathcal{O}_{r,k}\left(x^{\frac{1}{2}+\epsilon}\right).$